# Kitashirakawa Oiwakecho, Sakyo-ku, Kyoto 606-8502, Japan

### v1

@@ 3 Connected wedge theorem in AdS3/CFT2
The AdS/CFT correspondence is a duality between the asymptotically anti-de Sitter (AdS) bulk and a conformal field theory (CFT) on the boundary. In the semiclassical limit, this duality suggests that a quantity defined in the holographic CFT can be computed by a geometric quantity in the bulk and vice versa. For instance, the entanglement entropy of a state in a holographic CFT can be calculated as the area of the associated extremal surface. This duality also applies to the dynamics, revealing an illuminating interplay between the quantum operations and causality. In this section, we review one particular application, known as the connected wedge theorem May:2019yxi .
@@ 3.1 Holographic scattering from the asymptotic boundary
Figure 2: Two realizations of a quantum task. a) A local implementation: Applying a unitary gate $U$ between Alice and Bob’s qubits via direct scattering in the bulk. b) An equivalent, nonlocal realization: the task can be achieved on the boundary/screen without direct contact using an entangled resource state between Alice and Bob. White boxes represent some local operations. This is known as a nonlocal quantum computation.
In this subsection, we consider an asymptotic quantum task via holographic scattering. Let us consider a 2-to-2 scattering in AdS3/CFT2. The input and output points are placed on the boundary so that the causal future of the input points and the causal past of the output points have an intersection in the bulk. For the connected wedge theorem, we perform a quantum task at the scattering region. For example, qubits can be sent from each input point, Alice and Bob, and a unitary gate acts on them to process these qubits, then Alice and Bob receive their qubits at each output point (and perform local operations). This protocol accomplishes the quantum task through the direct scattering as shown in Figure 2a. Contrary to the bulk picture, there are cases where no direct scattering is possible on the boundary. Thus, we need to find a way to perform the same task nonlocally. Such a task is known as a nonlocal quantum computation. In general, entanglement between Alice and Bob is required to accomplish the task nonlocally (Figure 2b). By considering a particular task called the B84, it can be proven that to accomplish the task there must be a finite correlation between Alice and Bob. The relevant correlation must be present after than each input but before each qubit is sent to each other (Figure 2b). This implies that there must be a finite mutual information between each decision region $\mathcal{R}_{1,2}$, defined as the intersection of the future domain of dependence of each input and the past domains of dependence of two outputs:
@@
where $S(\mathcal{R})$ is the entanglement entropy between $\mathcal{R}$ and its complement. Note that the definition of the decision regions heavily relies on the causal structure on the boundary $\mathcal{S}$. To accomplish the B84 task $n$ times in parallel with high probability, one can show that the mutual information must be $I(\mathcal{R}_{1}:\mathcal{R}_{2})\gtrsim n$. By taking $n$ as large as $O(1/G_{N}^{a})$ with $a<1$ to avoid backreaction, one can argue that the mutual information must be $O(1/G_{N})$ due to the large-$N$ hierarchy between the area term and the quantum correction. In summary, the connected wedge theorem is given as follows:
@@
then the decision regions $\mathcal{R}_{1,2}=J^{+}(c_{1,2})|_{\mathcal{S}}\cap J^{-}(r_{1})|_{\mathcal{S}}\cap J^{-}(r_{2})|_{\mathcal{S}}$ on the boundary must have a connected entanglement wedge, implied from
@@
Note that the domains of dependence on the boundary are determined by the standard causality, namely, the conformally flat metric. Thanks to the Gao-Wald theorem, there is no causality discrepancy between the bulk and the boundary Omiya:2021olc .
This discussion is based on the position-based cryptography and does not rely on the detail of the theory Buhrman_2014 . Thus, it should work in any holographic spacetime. In this paper, we use this connected wedge theorem as a guiding principle to constrain or check various proposals related to the holographic duality. One application in the previous study will be reviewed in the next subsection.
@@ 3.2 Holographic scattering from a non-asymptotic boundary
The authors of Mori:2023swn , including one of us, extended the connected wedge theorem to a braneworld or cutoff surface in an asymptotically AdS3 spacetime. In these setups, the holographic screen $\mathcal{S}$ is located at somewhere other than the asymptotic boundary. It turns out that the causality based on the induced metric on $\mathcal{S}$ leads to an apparent violation of the connected wedge theorem. The resolution presented in the work is to fill behind the hypersurface with a fictitious asymptotically AdS space, [^3] In the work Mori:2023swn , the focus was a braneworld/cutoff AdS so just extending the original spacetime beyond the hypersurface was sufficient to resolve the puzzle. In general, one can glue an arbitrary fictitious spacetime with a fictitious asymptotic boundary as long as it satisfies the Israel junction condition Israel:1966rt to be a smooth spacetime satisfying the Einstein equation. and extend the scattering trajectories to the fictitious asymptotic boundary to define fictitious input and output points $\tilde{c}_{1,2}\in J^{-}(c_{1,2})$, $\tilde{r}_{1,2}\in J^{+}(r_{1,2})$. See Figure 3 for its illustration. The authors have shown that the boundary domains of dependence defined from the induced causality, denoted by $\hat{J}^{\pm}$, align with the connected wedge theorem. After all, Mori:2023swn proposes the following refined connected wedge theorem:
@@
Unlike the cases considered in Mori:2023swn , where holographic entanglement entropy prescription is examined in some explicit models on the brane or a cutoff surface, so far there is no explicit model of static patch holography for dS3. Hence the holographic entanglement entropy prescription is not verified from the dual quantum theory. However, by using the fact that various proposals reduce to a single prescription in static patch (Appendix A.3), we assume the prescription and ask what causality on the screen is consistent with the dS holography. As the connected wedge theorem has its origin in quantum information, it should be true regardless of the dual UV theory. Thus, the theorem offers a nontrivial criterion on causality and a consistency check with the holographic entanglement entropy prescription.
@@ 4 Static patch holography
In parallel to the development of holography and quantum gravity, mainly under the scope of AdS spacetime, there has been strong evidence that our universe has gone through two periods of accelerated expansion, that are the cosmic inflation in the early universe and the present period SupernovaSearchTeam:1998fmf ; SupernovaCosmologyProject:1998vns ; Planck:2015fie . An extensive understanding of the early universe asks for a quantum gravity description, and extending the great progress made in AdS/CFT to expanding universes is one of the major challenges in modern high-energy physics.
@@
where $\sigma\in(-\pi/2,\pi/2)$, $\theta\in[0,\pi]$, and $\varphi\in[0,2\pi)$. Here the radius of curvature $l_{dS}$ has been set to $1$. Future and past null infinities $\mathcal{I}^{\pm}$ are located at $\sigma=\pm\pi/2$, respectively. Once an observer is defined as a causal worldline in spacetime, de Sitter spacetime has the particularity to possess observer-dependent horizons. Indeed, the region that an observer can send and receive signals never covers the whole universe. Such a region is often called the static patch – or causal region – of the observer and it is bounded by a cosmological horizon.
Contrary to black hole horizons, cosmological horizons depend on the worldline of the observer, as they are constructed as the union of the causal past and causal future of the future and past endpoints of the worldline. The horizon can be made manifest by writing the dS metric in static coordinates,
@@
Following Susskind:2021omt ; Franken:2023pni , let us make the static patch holography proposal more precise. Consider an arbitrary observer in asymptotically dS spacetime. Achronal slices of the bulk are denoted by $\Sigma$. In the spirit of Franken:2023pni , we define the holographic screen associated with an observer as follows.
@@
The holographic screen $\mathcal{S}$ associated with an observer is the codimension-one timelike boundary of a region in which all closed codimension-two surfaces for which null geodesics orthogonal to the surface and directed towards the observer’s worldline are of nonpositive expansion. When referring to a holographic screen without mentioning an observer, $\mathcal{S}$ may be the union of holographic screens associated with different observers.
In pure dS, $\mathcal{S}$ is any timelike codimension-one surface inside the cosmological horizon. When it is located at a distance of order $O(_{N})$ away from the horizon, it is called a stretched horizon. Additionally, our definition includes holographic screens associated with finite-lifetime observers. Consider an observer living on $\sigma\in[\sigma_{a},\sigma_{b}]$, with associated worldline $\gamma_{M}$. One can construct its associated holographic screen by anchoring a codimension-two timelike screen to the endpoints of the observer worldline $\partial\gamma_{M}$ and continuing this screen along the worldline of an infinite-lifetime observer whose wordline contains $\gamma_{M}$. [^5] We thank Cynthia Keeler for discussions on this point.
@@
The semiclassical gravity description of the region inside the screen $\mathcal{S}$ is dual to a quantum theory defined on $\mathcal{S}$ [^6] Based on earlier works Susskind:dSentropy ; Shaghoulian:2021cef ; Shaghoulian:2022fop , it was argued by the authors of Franken:2023pni , including one of us, that two stretched horizons associated with two antipodal observers can not only encode their associated interior but also the bulk region separating them. See Appendix A.1 for more details.
This conjecture was initially motivated by the covariant entropy bound Susskind:2021omt ; Bousso:1999xy ; Bousso:2002ju and then developed in Susskind:2021dfc ; Susskind:2021esx ; Lin:2022nss ; Susskind:2023hnj . See Shyam:2021ciy ; Lewkowycz:2019xse ; Coleman:2021nor ; Banihashemi:2022htw ; Banihashemi:2022jys for additional supportive evidence. Moreover, a recent proposal of Conjecture 4.1 for asymptotically dS3 identifies $\mathcal{S}$ with either the worldline of the two antipodal observers or the two stretched horizons Susskind:2021esx ; Lin:2022nss ; Rahman:2022jsf ; Goel:2023svz ; Narovlansky:2023lfz ; Verlinde:2024znh ; Verlinde:2024zrh ; Blommaert:2023opb ; Blommaert:2023wad ; Rahman:2024iiu . Following a dimensional reduction, the dual theory is conjectured to be the double-scaled Sachdev-Ye-Kitaev (DSSYK) model, with a very encouraging number of nontrivial matches, including correlation functions, the partition function, and quasinormal modes. Conjecture 4.1 will be one of the main assumptions in our proof of the connected wedge theorem in the static patch (Theorem 6.1). By abuse of notation, we use the same name and notation, a screen $\mathcal{S}$, to represent both a timelike hypersurface $\mathcal{S}$ and the holographic quantum degrees of freedom on it $\mathcal{H}_{\mathcal{S}}$.

### v3

@@ 3 Connected wedge theorem in AdS3/CFT2
The AdS/CFT correspondence is a duality between the asymptotically anti-de Sitter (AdS) bulk and a conformal field theory (CFT) on the boundary. In the semiclassical limit, this duality suggests that a quantity defined in the holographic CFT can be computed by a geometric quantity in the bulk and vice versa. For instance, the entanglement entropy of a state in a holographic CFT can be calculated as the area of the associated extremal surface. This duality also applies to the dynamics, revealing an illuminating interplay between quantum operations and causality. In this section, we review one particular application, known as the connected wedge theorem May:2019yxi .
@@ 3.1 Holographic scattering from the asymptotic boundary
Figure 3: Two realizations of a quantum task. a) A local implementation: Applying a unitary gate $U$ between Alice and Bob’s qubits via a direct scattering in the bulk. b) An equivalent, nonlocal realization: the task can be achieved on the boundary/screen without direct contact using an entangled resource state between Alice and Bob. White boxes represent some local operations. This is known as a nonlocal quantum computation.
In this subsection, we consider an asymptotic quantum task via holographic scattering in AdS3/CFT2. Let Alice and Bob each have a qubit. They input each qubit at their respective input locations $c_{1},c_{2}$, then process them through a holographic evolution, and receive their qubits back again at their respective output locations $r_{1},r_{2}$. For the connected wedge theorem, we first consider performing a certain quantum task in the bulk through a direct scattering. The input and output points are placed on the boundary so that the causal future of the input points and the causal past of the output points have an intersection in the bulk, which is the scattering region. We configure the holographic evolution such that a unitary gate acts locally on the qubits in the scattering region to accomplish the task. This protocol accomplishes the quantum task through a direct scattering as shown in Figure 3a. [^3] Note that in principle, any unitary process can be implemented in a holographic setting. Now, let us turn to the boundary perspective. Since we work in holography, the task must be also achievable on the boundary, however, there are cases where no direct scattering is possible on the boundary. Thus, we need to find a way to perform the same task nonlocally. Such a task is known as a nonlocal quantum computation PhysRevLett.90.010402 ; Buhrman_2014 . In general, entanglement between Alice and Bob is required to accomplish the task nonlocally (Figure 3b). By considering a particular task called the B84, it can be proven that to accomplish the task there must be a finite correlation between Alice and Bob. The relevant correlation must be available after the input is received and in the past of both outputs (Figure 3b). This implies that there must be a finite mutual information between each decision region $\mathcal{R}_{1,2}$, defined as the intersection of the future domain of dependence of each input and the past domains of dependence of two outputs. The mutual information of these two regions is defined as
@@
where $S(\mathcal{R})$ is the entanglement entropy between $\mathcal{R}$ and its complement. Note that the definition of the decision regions heavily relies on the causal structure on the boundary $\mathcal{S}$. To accomplish the B84 task $n$ times in parallel with high probability, one can show that the mutual information must be $I(\mathcal{R}_{1}:\mathcal{R}_{2})\gtrsim n$. By taking $n$ any function as large as $o(1/G_{N})$ to avoid backreaction, one can argue that the mutual information must be $O(1/G_{N})$ May:2021nrl . [^4] The notation $f(x)=o(g(x))$ as $x\rightarrow\infty$ means that for any constant $\epsilon>0$ there exist $x_{0}$ such that $|f(x)|\leq\epsilon g(x)$ for $x\geq x_{0}$. Note that this does not rely on the large-$N$ separation between the area term and the quantum correction of the mutual information. [^5] This point is important, since later the dual quantum theory remains unknown so holographic entanglement entropy may not have a nice split into the area term of order $1/G_{N}$ (or possibly higher order) and quantum corrections of order unity. There is possibly a term growing with a rate that is sub-linear but larger than any $1/G_{N}^{a}$ with $a<1$. Even if this is the case, the argument in the main body says the mutual information must be at least as large as $1/G_{N}$, implying a geometric correlation from the area terms. We thank Alex May for pointing out a loophole in the first version of this paper, and explaining this to us.
By taking $n$ as large as $O(1/G_{N}^{a})$ with $a<1$ to avoid backreaction, one can argue that the mutual information must be $O(1/G_{N})$ due to the large-$N$ hierarchy between the area term and the quantum correction. In summary, the connected wedge theorem is given as follows:
@@
then the decision regions $\mathcal{R}_{1,2}=J^{+}(c_{1,2})|_{\mathcal{S}}\cap J^{-}(r_{1})|_{\mathcal{S}}\cap J^{-}(r_{2})|_{\mathcal{S}}$ on the boundary must have a connected entanglement wedge, equivalently
@@
This discussion is based on quantum information and does not rely on the details of the theory. [^6] One caveat is that we ignore the complementary regions of the decision regions. As pointed out in May:2019odp and the erratum of May:2019yxi , GHZ-type resource states among the complementary regions and $\mathcal{R}_{1}\cup\mathcal{R}_{2}$ can offer a protocol that does not need a finite mutual information between $\mathcal{R}_{1}$ and $\mathcal{R}_{2}$. This is however unlikely the case for holographic scattering as we do not expect a large amount of the GHZ-type entanglement in holography Susskind:2014yaa ; Nezami:2016zni ; Mori:2024gwe . Thus, it should work in any holographic spacetime. In this paper, we use this connected wedge theorem as a guiding principle to constrain or check various proposals related to the holographic duality. One application in the previous study will be reviewed in the next subsection.
@@ 3.2 Holographic scattering from a non-asymptotic boundary
The authors of Mori:2023swn , including one of us, extended the connected wedge theorem to a braneworld or cutoff surface in an asymptotically AdS3 spacetime. In these setups, the holographic screen $\mathcal{S}$ is located somewhere other than the asymptotic boundary. It turns out that the causality based on the induced metric on $\mathcal{S}$ leads to an apparent violation of the connected wedge theorem. The resolution presented in the work is to fill behind the hypersurface with a fictitious asymptotically AdS space, [^7] In the work Mori:2023swn , the focus was a braneworld/cutoff AdS so just extending the original spacetime beyond the hypersurface was sufficient to resolve the puzzle. In general, one can glue an arbitrary fictitious spacetime with a fictitious asymptotic boundary as long as it satisfies the Israel junction condition Israel:1966rt to be a smooth spacetime satisfying the Einstein equation. and extend the scattering trajectories to the fictitious asymptotic boundary to define fictitious input and output points $\tilde{c}_{1,2}\in J^{-}(c_{1,2})$, $\tilde{r}_{1,2}\in J^{+}(r_{1,2})$. See Figure 4 for its illustration. The authors have shown that the boundary domains of dependence defined from the induced causality, denoted by $\hat{J}^{\pm}$, align with the connected wedge theorem. After all, Mori:2023swn proposes the following refined connected wedge theorem:
@@
Unlike the cases considered in Mori:2023swn , where holographic entanglement entropy prescription is examined in some explicit models on the brane or a cutoff surface, so far there is no explicit model of static patch holography for dS3. Hence the holographic entanglement entropy prescription is not verified from the dual quantum theory. However, by using the fact that various proposals reduce to a single prescription in a static patch (Appendix B.3), we assume the prescription and ask what causality on the screen is consistent with the dS holography. As the connected wedge theorem has its origin in quantum information, it should be true regardless of the dual UV theory. Thus, the theorem offers a nontrivial criterion on causality and a consistency check with the holographic entanglement entropy prescription.
@@ 4 Static patch holography
where $\sigma\in(-\pi/2,\pi/2)$, $\theta\in[0,\pi]$, and $\varphi\in[0,2\pi)$. See Figure 5 for the associated Penrose diagram.
Figure 5: Penrose diagram of de Sitter space. Two horizontal lines are the past and future null infinities $\mathcal{I}^{\pm}$. The shaded region is the causal region – static patch – of a static observer located at $\theta=\pi$ (south pole).
Here the radius of curvature $l_{dS}$ has been set to $1$. Future and past null infinities $\mathcal{I}^{\pm}$ are located at $\sigma=\pm\pi/2$, respectively. Once an observer is defined as a causal worldline in spacetime, de Sitter spacetime has the particularity to possess observer-dependent horizons. Indeed, the region where an observer can send and receive signals never covers the whole universe. Such a region is often called the static patch – or causal region – of the observer and it is bounded by a cosmological horizon. [^9] As mentioned in footnote 2, the static patch in this paper is not necessarily described by a time-independent metric.
Contrary to black hole horizons, cosmological horizons depend on the worldline of the observer, as they are constructed as the union of the causal past and causal future of the future and past endpoints of the worldline. The observer’s and antipodal observer’s worldlines are given by $\theta=\pi$ (south pole) and $\theta=0$ (north pole) in the conformal coordinates, respectively. The cosmological horizon can be made manifest by writing the dS metric in static coordinates,
@@
Following Susskind:2021omt ; Franken:2023pni , let us make the static patch holography proposal more precise. Consider an arbitrary observer in asymptotically dS spacetime. Achronal slices of the bulk are denoted by $\Sigma$. To encode bulk dynamics holographically, a holographic screen should be a timelike, codimension-one hypersurface. Furthermore, we expect that the (code subspace) algebra of the holographic screen should be equal to that of the observer’s worldline $\zeta$. Since the timelike tube theorem Borchers1961 ; osti_4665531 ; Strohmaier:2023opz states that the algebra of the observer’s worldline is equivalent to that of operators in its timelike envelope $E(\zeta)$ Witten:2023xze , the domain of dependence of the holographic screen should equal $E(\zeta)$. This means $\partial\zeta$ should coincide with the future and past timelike edge of the holographic screen. [^10] Physically, this means no signal from the observer travels beyond the holographic screen without crossing it. Conversely, any signal sent from a point beyond the screen cannot be received by the observer without crossing the screen. Additionally, the holographic screen should be convex to avoid any subtleties regarding holographic entanglement entropy prescription and induced causality Mori:2023swn . [^11] This condition could possibly be removed by considering a generalized entanglement wedge Bousso:2022hlz . After these considerations, we define the holographic screen in static patch holography as follows:
Definition 4.1 (Holographic screen in static patch of asymptotically dS).
The holographic screen $\mathcal{S}$ associated with an observer is a codimension-one convex timelike hypersurface in the static patch, anchored to the observer’s worldline endpoints.
Remark 1.
We can generalize this notion of the holographic screen to more general spacetimes such as a Friedmann-Lemaître-Robertson-Walker spacetime by explicitly requiring non-positive expansion toward the observer’s worldline to ensure the Bousso bound Bousso:1999xy as follows:
@@
The holographic screen $\mathcal{S}$ associated with an observer is the codimension-one convex timelike boundary of a region in which all closed codimension-two surfaces for which null geodesics orthogonal to the surface and directed towards the observer’s worldline are of non-positive expansion. When referring to a holographic screen without mentioning an observer, $\mathcal{S}$ may be the union of holographic screens associated with different observers.
In pure dS, $\mathcal{S}$ is any convex timelike codimension-one surface inside the cosmological horizon. When it is located at a distance of order $O(G_{N})$ away from the horizon, it is called a stretched horizon. Note that our definition naturally generalizes to holographic screens associated with finite-lifetime observers. [^12] We thank Cynthia Keeler for discussions on this point.
@@
The semiclassical gravity description of the region inside the screen $\mathcal{S}$ is dual to a quantum theory defined on $\mathcal{S}$ [^13] Based on earlier works Susskind:2021esx ; Shaghoulian:2021cef ; Shaghoulian:2022fop , it was argued by the authors of Franken:2023pni , including one of us, that two stretched horizons associated with two antipodal observers can not only encode their associated interior but also the bulk region separating them. See Appendix B.1 for more details.
This conjecture was initially motivated by the covariant entropy bound Susskind:2021omt ; Bousso:1999xy ; Bousso:2002ju and then developed in Susskind:2021dfc ; Susskind:2021esx ; Lin:2022nss ; Susskind:2023hnj . See Shyam:2021ciy ; Lewkowycz:2019xse ; Coleman:2021nor ; Banihashemi:2022htw ; Banihashemi:2022jys for additional supportive evidence. Moreover, a recent proposal of Conjecture 4.1 for asymptotically dS3 identifies $\mathcal{S}$ with either the worldline of the two antipodal observers or the two stretched horizons Susskind:2021esx ; Lin:2022nss ; Rahman:2022jsf ; Goel:2023svz ; Narovlansky:2023lfz ; Verlinde:2024znh ; Verlinde:2024zrh ; Blommaert:2023opb ; Blommaert:2023wad ; Rahman:2024iiu ; Milekhin:2023bjv ; Xu:2024hoc ; Milekhin:2024vbb . Following a dimensional reduction, the dual theory is conjectured to be the double-scaled Sachdev-Ye-Kitaev (DSSYK) model, with a number of nontrivial matches, including correlation functions, the partition function, quasinormal modes, hyperfast scrambling and the algebra of observables. Conjecture 4.1 will be one of the main assumptions in our proof of the connected wedge theorem in the static patch (Theorem 6.1).
To compute holographic entanglement entropy of subsystems of spatial slices of the holographic screen, we will consider the bilayer proposal Shaghoulian:2021cef ; Franken:2023pni . A detailed introduction to this prescription is provided in Appendix B. Consider a spacelike subsystem of the screen, $A\in\mathcal{S}|_{\Sigma}$. Its holographic entanglement entropy in the proposal is given by the sum of the areas of the two extremal surfaces homologous to $A$ on both sides of the screen, $D$ and $D_{E}$, where $D$ denotes the codimension-zero interior of the screen, and $D_{E}$ denotes the codimension-zero region sandwiched by the screen and the complementary one associated with an antipodal observer. For their precise definition, see Appendix B and Fig. 12. (Note that $D$ is denoted as $D_{R}$ in the appendix.) The holographic entanglement entropy is proposed as
$S(A)=\frac{1}{4G_{N}}\left[\mathrm{area}(\gamma_{e}(A;D))+\mathrm{area}(\gamma_{e}(A;D_{E}))\right]+O(1),$ (18)
where $\gamma_{e}(A;D)$ and $\gamma_{e}(A;D_{E})$ are the minimal extremal surfaces homologous to $A$ in $D$ and $D_{E}$, respectively.
We show in Appendix B.3 that analogous to the AdS case, $\gamma_{e}(A;D)$ can be found by a maximin procedure, which amounts to finding the minimal area surface on any spatial slice $\Sigma$ such that $\partial\Sigma=\mathcal{S}|_{\Sigma}$, and choosing the maximum-area surface among them:
$\mathrm{area}(\gamma_{e}(A;D))=\max_{\Sigma}\left(\min_{\gamma\in\Sigma}\left(\mathrm{area}(\gamma)\right)\right).$ (19)
In contrast, $\gamma_{e}(A;D_{E})$ depends on which extremal surface prescription we use, as mentioned in Appendix B.2. In particular, extremal surfaces are not always maximin surfaces in $D_{E}$, and conversely. Nevertheless, it does not matter for the discussion of the connected wedge theorem as explained now. The homology condition in $D_{E}$ suggests that any subsystem of the screen is not homologous to its complementary regions on the screen, due to the presence of another boundary. Thus, given two spacelike subsystems $A$ and $B$ of the screen, $\gamma_{e}(A\cup B;D_{E})$ is necessarily equal to $\gamma_{e}(A;D_{E})\cup\gamma_{e}(A;D_{E})$. It follows that the contributions from $D_{E}$ cancel out in the mutual information. Finally, the mutual information is purely given in terms of the area in the interior as
$I(A:B)=\frac{1}{4G_{N}}\left[\mathrm{area}(\gamma_{e}(A;D))+\mathrm{area}(\gamma_{e}(B;D))-\mathrm{area}(\gamma_{e}(A\cup B;D))\right].$ (20)